Possible Survey Questions Suggestions – Ask about five from each group (or similar questions). Whichever of these you choose, provide a 5-item response spectrum: 1 = Strongly Disagree, 2 = Disagree, 3 = Neutral, no opinion, 4 = Agree, 5 = Strongly Agree.

Individualist – Communitarian Axis

1. In a fair system, people with more ability should have the opportunity to rise as far as they can.

2. Collaboration with others is necessary in today’s world and leads to results superior to those that could be obtained by working alone.

3. I am happiest when I’m in a group of like-minded friends.

4. It is right that life tends to sort out those who try harder from those who don’t.

5. I don’t tend to join clubs or make associations readily.

6. Preserving citizens’ freedom is more important than other goals such as minimizing terrorism or violent crime.

7. Most people make friends only because friends are useful to them.

Hierarchical – Egalitarian Axis

1. What the current U.S. needs is a “fairness revolution” to make the distribution of wealth more equal.

2. Important decisions about our future should be made by experts rather than the general public.

3. I have a high degree of respect for law enforcement personnel and the military.

4. I prefer always having clear instructions from teachers or supervisors about what to do.

5. All workers in a firm should have input into important business decisions that can affect their jobs.

6. An intact family is the basis of a functioning society.

7. Minimum wage laws are an appropriate way to guarantee everyone a decent living.